# Parametric Instabilities of Correlated Quantum Matter

### v1

@@ V Discussion and outlook
Parametric driving of non-trivial low-energy collective excitations is a new and exciting avenue for controlling, engineering, and probing many-body correlated states. The driven dynamics proposed here are enabled by recent experimental advances, the large degree of tunability, and the rich phase diagram that new correlated electronic platforms allow. We find that modulating microscopic parameters at the right frequency, allows one to approach phenomena that are otherwise practically inaccessible in the solid-state context.
@@
The measurable signature of the parametric drive, i.e., the amplitude of its response, may be system-specific, and depends on the detailed character of the low-energy ordered state. In the excitonic double layer example (Sec. IV.1), interlayer tunneling spectroscopy reveals clear signatures of the parametrically driven state. Systems with richer structure, e.g., the flat-band “pseudo-TBG” model we explore in Sec. IV.2, has more detection possibilities, such as time-varying charge-density ordering and modulated optical responses. In all cases, however, the parametric response is always sub-harmonic with respect to the driving frequency. All of our examples have $\omega_{d}$ corresponding to twice the frequency of the targeted collective mode, and thus the parametric response would prominently appear at a frequency (or energy corresponding to) $\omega_{d}/2$.
@@
A key result addresses the strength of the induced parametric response. We find that the parametric susceptibility is intimately connected to the nature of the vacuum state and its dependence on microscopic parameters. As such, parametric driving emerges as a novel probe of the fidelity susceptibility in the context of solid-state materials and devices. This quantity generically gets enhanced and diverges in the vicinity of quantum phase transitions, where the nature of the vacuum ground-state changes dramatically. Therefore, mapping the readily-implemented parametric driving response across swaths of phase space unlocks a new and complimentary way to examine quantum properties in correlated materials, detect quantum fluctuations, and discover hidden phase transitions, which may be otherwise hard to reach.
@@
The non-trivial connection between the vacuum fidelity susceptibility and the induced parametric drive was revealed through a framework that included a quantum theory of the collective many-body fluctuations. Importantly, our analysis emphasized the dynamical properties of these collective modes, and not only their energetics (i.e., their spectrum). The emergent conjugate-canonical-pair structure of the collective coordinates played a key role, and our approach enables straightforward identification of such dynamical-variable pairs. This is achieved by a special kind of partial bosonic Bogoliubov transformation, which isolates the parametric two-boson response of a system to certain changes of its parameters.
@@
In our analysis, squeezing emerged as a main organizing principle in parametric driving of electronically ordered materials. The two quadratures of the fluctuation, analogous to “position” and “momentum” (e.g., spin fluctuations in perpendicular directions), are to be treated on equal footing from the outset. The low-energy Hamiltonian, however, may then differentiate between the quadratures, and produce squeezing of the vacuum fluctuations. To induce parametric driving, a modulation must alter the squeezing, which corresponds to inducing an anomalous (i.e., two-boson) driving terms. Indeed, the more squeezing the reference ordered state has, the stronger is its parametric response (see, e.g., Eq. (38)). Intuitively, a squeezed vacuum state simply hosts a wider range of bosonic fluctuations.
@@
The case studies presented in Sec. IV also showed the role quantum geometry has as a driven parameter with associated parametric instabilities. In strongly interacting, narrow bandwidth, electronic systems, quantum geometry is crucial. Indeed, the competition between the geometry-associated length scale and the range of intrinsic electronic interactions is key in determining the parametric susceptibility with respect to small changes in the quantum geometry (as evident in Fig. 5b).
@@
Our framework concentrated on periodic modulations, but could also be applied to quantum parameter quenches. In a small quench, the reference electronic order remains intact, but it could excite the low-energy collective degrees of freedom. Quantifying this response could use the method summarized in Eq. (24). Inevitably, if the parametric susceptibility associated with the quench is substantial, a non-negligible amount of bosonic collective excitations will emerge. Such excitations, recently probed in correlated optically-pumped moiré systems, may be long-lived and have intriguing consequences in the context of characterizing the underlying electronic order [135, 136]. Also, one may ask how much work does the quench produce through exciting low-lying collective modes. Once again, the fidelity of the fluctuation vacuum is expected to play the key quantitative role. Interestingly, probing the statistics of the performed work probes the corresponding Loschmidt echo [137, 138], which has thus far remained out of experimental reach in the context of electronic condensed matter systems.
@@
An exciting future direction would be harnessing the parametric driving described here in quantum sensors and quantum information processing devices, specifically in the context of quantum-limited parametric amplification [139, 140, 141, 142, 143]. For instance, An excitonic strongly correlated double layer (see Sec. IV.1) could perhaps amplify a quantum signal fed into one of its gate voltages. A parametric instability may be triggered, exciting two Goldstone bosons. This will result in an oscillating interlayer electric field, which may be extracted as the amplified signal. This scheme provides phase-sensitive degenerate parametric amplification, where one of the quadratures of the signal is amplified and the other attenuated. A non-degenerate amplification setup can be explored when (i) several low-lying excitation modes are available, and (ii) joint parametric terms may be driven, e.g., $\sim a_{1}^{\dagger}a_{2}^{\dagger}$. (This is the case for our studied flat-band model, see lower right panel of Fig. 1c.)
@@
Finally, we speculate that effortless manipulation of the emergent bosonic collective excitations in the correlated electronic systems may have broad implications. These excitations may be coupled to other bosonic modes in various settings, such as mechanical oscillators [144, 145], microwave resonators [146, 147], superconducting qubits [147, 148], and optical cavities [149, 150]. Integrating solid-state tunable correlated electronic materials with these more established traditional quantum platforms holds great promise for applications of hybrid quantum systems [151, 152], e.g., quantum transduction [153, 154, 155, 156] and information storage [157, 158]. Controllably coupling the collective excitations explored here to other bosonic modes remains a fascinating topic for future work.
@@ Appendix F Flat band model
The parameter $\zeta$ controls the quantum metric of the projected flat bands. The Hamiltonian Eq. (52a) has a total of eight bands: four valence bands sitting at t-t, and four conduction bands at +t+t (when $\Delta_{\sigma}=0$). Our interest will lie in one set of these isolated bands, e.g., the valence bands. Within such set, one may define a good quantum number which further classifies the bands, ${\cal C}=\sigma_{z}\tau_{z}$.
@@ F.2 Fluctuation Lagrangian
Performing the crude lowest-mode approximation, we assign coordinates to the different generators in the following manner, $\tau_{z}\to{\mathbb{X}_{\bf Q}^{\rm G}}$, $\sigma_{x}\tau_{y}\to{\mathbb{P}_{\bf Q}^{\rm G}}$, $\sigma_{z}\to{\mathbb{X}_{\bf Q}^{\rm K}}$, $\sigma_{y}\tau_{x}\to{\mathbb{P}_{\bf Q}^{\rm K}}$. We are going to focus below on the ${\bf Q}\to 0$ contribution of these terms. One interesting “casualty” of this approximation is embodied by the last two lines in the expression above. Specifically, this term captures a coupling at larger moment between the gapless Goldstone mode and the so-called “K” mode. Within this approximation, we can finally find the contribution

### v2

@@ V Discussion and outlook
Parametric driving of non-trivial low-energy collective excitations is a new and exciting avenue for controlling, engineering, and probing many-body correlated states. The driven dynamics proposed here are enabled by recent experimental advances, the large degree of tunability, and the rich phase diagram that new correlated electronic platforms allow. We find that modulating the right microscopic parameters at an appropriate resonant frequency, allows one to approach phenomena that are otherwise practically inaccessible in the solid-state context.
@@
The measurable signature of the parametric drive may be system-specific, and depends on the detailed character of the low-energy ordered state. In the excitonic double layer example (Sec. IV.1), interlayer tunneling spectroscopy reveals clear signatures of the parametrically driven state. Systems with richer structure, e.g., the flat-band TBG-analog model we explore in Sec. IV.2, have more detection possibilities, such as time-varying charge-density ordering and modulated optical responses. In all cases, however, the parametric response is always sub-harmonic with respect to the driving frequency. All of our examples have $\omega_{d}$ corresponding to twice the frequency of the targeted collective mode, and thus the parametric response would prominently appear at a frequency (or energy corresponding to) $\omega_{d}/2$.
@@
A key result addresses the strength of the induced parametric response. We find that the parametric susceptibility is intimately connected to the nature of the vacuum state and the dependence of its properties on microscopic parameters. As such, parametric driving emerges as a novel probe of the fidelity susceptibility in the context of solid-state materials and devices. This susceptibility generically gets enhanced and diverges in the vicinity of quantum phase transitions, where the wavefunction of the vacuum ground-state changes dramatically. Therefore, mapping the readily-implemented parametric driving response across swaths of phase space unlocks a new and complimentary way to examine quantum properties in correlated materials, detect quantum fluctuations, and discover hidden phase transitions, which may be otherwise hard to reach.
@@
The non-trivial connection between the vacuum fidelity susceptibility and the induced parametric drive was revealed through a framework that included a quantum theory of the collective many-body fluctuations. Importantly, our analysis emphasized the dynamical properties of these collective modes, and not only their energetics (i.e., their spectrum). The emergent conjugate-canonical-pair structure of the collective coordinates [Eq. (16)] plays a key role, and our approach enables straightforward identification of such dynamical-variable pairs. This is achieved by a particular kind of partial bosonic Bogoliubov transformation, which isolates the parametric two-boson response of a system to certain changes of its parameters.
@@
In our analysis, squeezing emerged as a main organizing principle in parametric driving of electronically ordered materials. The two quadratures of the fluctuation, analogous to “position” and “momentum” (e.g., spin fluctuations in perpendicular directions), are to be treated on equal footing from the outset. The low-energy Hamiltonian, however [Eq. (17)], then differentiates between the quadratures, and produces squeezing of the vacuum fluctuations. To induce parametric driving, a modulation must alter the squeezing, which corresponds to inducing anomalous (i.e., two-boson) driving terms in the effective low-energy theory. Indeed, the more squeezing the reference ordered state has, the stronger is its parametric response (see, e.g., Eq. (40)). Intuitively, a squeezed vacuum state simply hosts a wider range of bosonic fluctuations.
@@
The case studies presented in Sec. IV showcase the role of quantum geometry as a potent driven parameter with associated parametric instabilities. In strongly interacting, narrow bandwidth, electronic systems, quantum geometry is crucial. Indeed, the competition between the geometry-associated length scale (e.g., $\zeta$) and the range of intrinsic electronic interactions is key in determining the parametric susceptibility with respect to small changes in the quantum geometry (as evident in Fig. 7b). These observations naturally motivate further studies into ways of manipulating and engineering quantum geometric quantities, e.g., Berry curvature and the quantum metric, in correlated quantum materials.
@@
Our framework concentrated on periodic modulations, but could also be applied to quantum parameter quenches. In a small quench, the reference electronic order remains intact, yet the low-energy collective degrees of freedom could be generically excited. Quantifying this response could use the method summarized in Eq. (26). Inevitably, if the parametric susceptibility associated with the quench is substantial, a non-negligible amount of bosonic collective excitations will emerge. Such excitations, recently probed in correlated optically-pumped moiré systems, may be long-lived and have intriguing consequences in the context of characterizing the underlying electronic order [141, 142]. Also, one may ask how much work does the quench produce through exciting low-lying collective modes. Once again, the fidelity of the fluctuation vacuum is expected to play the key quantitative role. Interestingly, probing the statistics of the performed work probes the corresponding Loschmidt echo [143, 144], which has thus far remained out of experimental reach in the context of electronic condensed matter systems.
@@
An exciting future direction would be harnessing the parametric driving described here in quantum sensors and quantum information processing devices, specifically in the context of quantum-limited parametric amplification [145, 146, 147, 148, 149]. For instance, An excitonic strongly correlated double layer (see Sec. IV.1) may be employed to amplify a photonic signal fed into one of its gate voltages. A parametric instability may be triggered, exciting two Goldstone bosons. This will result in an oscillating interlayer electric field, which may be extracted as the amplified signal. This scheme provides phase-sensitive degenerate parametric amplification, where one of the quadratures of the signal is amplified and the other attenuated. A non-degenerate amplification setup can be explored when (i) several low-lying excitation modes are available, and (ii) joint parametric terms may be driven, e.g., $\sim a_{1}^{\dagger}a_{2}^{\dagger}$. (This is the case for our studied flat-band model, see lower right panel of Fig. 2c.)
@@
Finally, we speculate that effortless manipulation of the emergent bosonic collective excitations in the correlated electronic systems may have broad implications. These excitations may be coupled to other bosonic modes in various settings, such as mechanical oscillators [150, 151], microwave resonators [152, 153], superconducting qubits [153, 154], and optical cavities [155, 156]. Integrating solid-state tunable correlated electronic materials with more established traditional quantum platforms holds great promise for applications of hybrid quantum systems [157, 158], e.g., quantum transduction [159, 160, 161, 162] and information storage [163, 164]. Controllably coupling the collective excitations explored here to other bosonic modes remains a fascinating topic for future work.
@@ Appendix F Flat band model
The parameter $\zeta$ controls the quantum metric of the projected flat bands. To see this, consider without loss of generality the $\sigma_{z}\tau_{z}=1$ set of bands. The Bloch eigenvector at momentum $\bf k$ of the valence band is given by
${\mathsf{u}}_{{\bf k}}=\frac{1}{\sqrt{2}}e^{i\alpha_{{\bf k}}\frac{\rho_{z}}{2}}\left(1,i\right)=\frac{1}{\sqrt{2}}\begin{pmatrix}e^{i\frac{\alpha_{{\bf k}}}{2}}\\ ie^{-i\frac{\alpha_{{\bf k}}}{2}}\end{pmatrix}$. (157)
Consider the Fubini-Study metric,
$g_{\mu\nu}={\rm Tr}\left[\partial_{\mu}P_{\bf k}\partial_{\nu}P_{\bf k}\right],$
with the band projections $P_{\bf k}^{\alpha}=|{\mathsf{u}}_{\bf k}^{\alpha}\rangle\langle{\mathsf{u}}_{\bf k}^{\alpha}|$. Straightforward calculation yields,
$g_{\mu\nu}\left({\bf k}\right)=\frac{\zeta^{2}}{4}\sin\left(k_{\mu}\right)\sin\left(k_{\nu}a\right).$ (158)
One may define the following quantum-geometric length scale,
$\ell_{{\rm geo.}}\equiv\sqrt{\int_{{\rm BZ}}\frac{d^{2}{\bf k}}{4\pi^{2}}{\rm Trg\left({\bf k}\right)}}=\frac{\zeta}{2},$ (159)
which corresponds to the minimal Wannier-function spread associated with these flat bands [165].
The Hamiltonian Eq. (53) has a total of eight bands: four valence bands sitting at t-t, and four conduction bands at +t+t (when $\Delta_{\sigma}=0$). Our interest will lie in one set of these isolated bands, e.g., the valence bands. Within such set, one may define a good quantum number which further classifies the bands, ${\cal C}=\sigma_{z}\tau_{z}$.
@@ F.2 Fluctuation Lagrangian
Performing the crude lowest-mode approximation, we assign coordinates to the different generators in the following manner, $\tau_{z}\to{\mathbb{X}_{\bf Q}^{\rm G}}$, $\sigma_{x}\tau_{y}\to{\mathbb{P}_{\bf Q}^{\rm G}}$, $\sigma_{z}\to{\mathbb{X}_{\bf Q}^{\rm K}}$, $\sigma_{y}\tau_{x}\to{\mathbb{P}_{\bf Q}^{\rm K}}$. We are going to focus below on the ${\bf Q}\to 0$ contribution of these terms. One interesting “casualty” of this approximation is embodied by the last two lines in the expression above. Specifically, this term captures a coupling at larger momenta between the gapless Goldstone mode and the so-called “K” mode, which may have some intriguing consequences beyond the scope of this work [166, 167, 78]. Within this approximation, we can finally find the contribution